OpenStudy (anonymous):

Solve. 12/=48/100 n =

OpenStudy (tkhunny):

Unfortunately, it cannot be solved as posted. Make a better effort to post legibly. Make ANY effort to show your own work.

OpenStudy (calculusxy):

Is it like: \[\frac{ 12 }{ n }=\frac{ 48 }{ 100 }\]

OpenStudy (anonymous):

Solve. 12/n = 48/100 n =

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Okay, now how about your efforts to solve?

OpenStudy (calculusxy):

So this means that you just have to multiply diagonally.

OpenStudy (anonymous):

i think it is 7

OpenStudy (calculusxy):

I don't think so...

OpenStudy (calculusxy):

You see that if you multiply diagonally you get:  \[12(100)=48(n)\]